# OtterROS: Picking and Programming an Uncrewed Surface Vessel for Experimental Field Robotics Research with ROS 2*

### v1

@@ I Introduction
Figure 1: The Otter USV during experiments in Lake Ontario. Encounters with pleasure craft are common, but occasionally we cross paths with history.
@@ II-B Middleware for USV Autonomy
First release
2007
2001
2017
Maintainer
LSTS
MRG and LAMSS
Open Robotics
Development origins
USVs
USVs
Generic
Published use [^3] A snapshot of search result count from March 2024 from Google scholar and GitHub for “DUNE Unified Navigation Environment”, “MOOS-IvP”, and “ROS 2” (and “ROS 2 USV”).
~500
~1,000
~30,000 (~100 with “USV”)
Documentation
Limited
MIT User Guide
ROS 2 Docs, YouTube, …
@@ IV-B Available Data (Publisher)
otter_gps
NavSatFix
Latitude, longitude, altitude
otter_gps_time
OtterTime
GPS time
otter_imu
Imu
Orientation, angular velocity
otter_status
OtterStatus
Motor RPM, temperature, power consumption
otter_cogsog
COGSOG
Course, speed over ground, local speed vector
@@ V-B Integration and Supporting Hardware
Power distribution
MR power board
1
Power input cable
MR bulkhead power harness
1
Power harness (internal)
45130-0203 to device
3
Network switch
Netgear GS308
1
Ethernet connectors
ROP-5SPFFH-TCU7001
3
Network cables
Cat6 (generic)
4
Payload computer
NVIDIA Jetson AGX Orin Developers Kit
1
USB bulkhead
UA-30PMFL-LC7B20
1
Switch and indicator
MPB16-CARE-6-JR-0V
1
@@ VI-B Test Conditions
  • air temperatures from –5 ${}^{\circ}$ to 25 ${}^{\circ}$;
@@ VI-C System Performance
Motor failure
Mechanical failure in a motor from unknown cause. Possible propeller strike or due to rapid control commands (to be determined).
Motor delay
The Otter USV has an input delay of approximately 2 s when starting from stationary. A small initial input may mitigate power on delay.
Communication range
Difficult to gauge when communications will be lost. Constant monitoring required if Otter is not on a return path.
Local network dropouts
In cold conditions, the Otter suffers from network dropouts spanning a few seconds. No solution yet.
@@
Motor blockage
Jammed motors may mean the USV will be unable to return to shore. Scout ahead before deploying boat and seek local experts to avoid any (possible) vegetation.
Ecosystem damage
USVs can access areas inaccessible by other powered craft, causing damage to mostly undisturbed environments. Researchers should connect with local ecosystem experts to determine if, where, and when it is appropriate for field testing.
@@
Legality
Unlike UAVs, we have not found any rules about USV usage. To avoid any conflict, consult with local authorities (e.g., Coast Guard, government agencies) before starting experiments.
Runtime
The Otter has more endurance than most laptop equipment and researchers. Access to power, shelter, and washrooms is critical for a full day of testing (especially in cold conditions).
Public interaction
Popular with public, so boaters and swimmers frequently approach the Otter. Limited signage provides no warning to nearby public and no built in audio connection to base station. Requires supervision and patience for interrupted experiments.
Sailing right of way
When far from shore and encountering another vessel, it is hard to gauge safe passage. Stopping the Otter provides the safest way to let the other vessel pass.

### v2

@@ I Introduction
Figure 1: The Otter USV (left) during experiments in Lake Ontario. Encounters with pleasure craft are common, but occasionally we cross paths with history.
@@ II-B Middleware for USV Autonomy
First release 2007 2001 2017
Maintainer LSTS MRG and LAMSS Open Robotics
Development origins USVs USVs Generic
Published use [^3] A snapshot of search result count from March 2024 from Google scholar and GitHub for “DUNE Unified Navigation Environment”, “MOOS-IvP”, and “ROS 2” (and “ROS 2 USV”). ~500 ~1,000 ~30,000 (~100 with “USV”)
Documentation Limited MIT User Guide ROS 2 Docs, YouTube, …
@@ IV-B Available Data (Publisher)
otter_gps NavSatFix Latitude, longitude, altitude
otter_gps_time OtterTime GPS time
otter_imu Imu Orientation, angular velocity
otter_status OtterStatus Motor RPM, temperature, power consumption
otter_cogsog COGSOG Course, speed over ground, local speed vector
@@ V-B Integration and Supporting Hardware
Power distribution MR power board 1
Power input cable MR bulkhead power harness 1
Power harness (internal) 45130-0203 to device 3
Network switch Netgear GS308 1
Ethernet connectors ROP-5SPFFH-TCU7001 3
Network cables Cat6 (generic) 4
Payload computer NVIDIA Jetson AGX Orin Developers Kit 1
USB bulkhead UA-30PMFL-LC7B20 1
Switch and indicator MPB16-CARE-6-JR-0V 1
@@ VI-B Test Conditions
  • air temperatures from –5 C to 25 C;
@@ VI-C System Performance
Motor failure Mechanical failure in a motor from unknown cause. Possible propeller strike or due to rapid control commands (to be determined).
Motor delay The Otter USV has an input delay of approximately 2 s when starting from stationary. A small initial input may mitigate power on delay.
Communication range Difficult to gauge when communications will be lost. Constant monitoring required if Otter is not on a return path.
Local network dropouts In cold conditions, the Otter suffers from network dropouts spanning a few seconds. No solution yet.
@@
Motor blockage Jammed motors may mean the USV will be unable to return to shore. Scout ahead before deploying boat and seek local experts to avoid any (possible) vegetation.
Ecosystem damage USVs can access areas inaccessible by other powered craft, causing damage to mostly undisturbed environments. Researchers should connect with local ecosystem experts to determine if, where, and when it is appropriate for field testing.
@@
Legality Unlike UAVs, we have not found any rules about USV usage. To avoid any conflict, consult with local authorities (e.g., Coast Guard, government agencies) before starting experiments.
Runtime The Otter has more endurance than most laptop equipment and researchers. Access to power, shelter, and washrooms is critical for a full day of testing (especially in cold conditions).
Public interaction Popular with public, so boaters and swimmers frequently approach the Otter. Limited signage provides no warning to nearby public and no built in audio connection to base station. Requires supervision and patience for interrupted experiments.
Sailing right of way When far from shore and encountering another vessel, it is hard to gauge safe passage. Stopping the Otter provides the safest way to let the other vessel pass.